# Locret: Enhancing Eviction in Long-Context LLM Inference with Trained Retaining Heads

### v1

@@ 4.3 Speed Test on Real Consumer-Grade Devices
Success
4.4 Orthogonality to Other Methods
We evaluate the combination of Locret with quantization and token merging to address Q3.
@@ 4.4 Orthogonality to Other Methods
KV cache quantization. According to Zhang et al. (2024d), eviction-based methods like H2O struggle with compatibility when combined with KV cache quantization. Quantization introduces significant disturbance in the estimation of heavy-hitters, leading to severe performance degradation. However, Locret is not affected by such issues and can be combined with quantization while maintaining most of its performance. Here, we compare the performance degradation caused by quantization on Locret with that of the full attention method using the same metrics. We use Quanto as the quantization backend and report the average L-Eval score with Llama-3.1-8B-instruct as the model backbone. Table 4.4 shows that the performance drop caused by quantization on Locret is only slightly higher than that observed with the full attention method, indicating that Locret is a quantization-friendly approach. More details of the experiment are provided in Appendix C.1.
@@
Token merging. As described in Section 2, token dropping can also be implemented through an attention pool. Attention pool-based methods (Xiao et al., 2024b; Cai et al., 2024; Mu et al., 2024; Munkhdalai et al., 2024) merge adjacent tokens or cache units into an attention pool, maintaining a static cache size. These methods are orthogonal to Locret , as the evicted tokens can be merged into a small cache pool and retained in GPU memory. We conduct the following experiment to demonstrate that Locret can serve as an effective plug-in scoring function within such frameworks, enhancing performance without increasing memory budget. We select LoCoCo (Cai et al., 2024) as a representative of the latest attention pool-based methods. LoCoCo maintains a cache set consisting of two parts: the heavy hitters and the convolved non-heavy hitters. During each chunked prefill step, LoCoCo first identifies a set of heavy hitters according to H2(Zhang et al., 2024e), then applies 1-D convolution to the non-heavy hitters to compress them into a static size. By replacing H2O’s heavy-hitter scoring function with Locret, we retain the cache units with high CIS and convolve the others. We compare this combination with standalone LoCoCo and Locret on L-Eval using the Llama-3.1-8B-instruct backbone and report the average score across all selected tasks. As shown in Table 4.4, Locret achieves a higher score than LoCoCo, and the combined algorithm outperforms both standalone methods. This suggests that Locret provides a more accurate scoring function compared to H2O, and the two methods complement each other, demonstrating their orthogonality. Further details of the experiment are provided in Appendix C.2.
5 Conclusion
We propose Locret, a lightweight training-based method that enables memory-efficient inference of long contexts on consumer-grade devices. Locret introduces a few additional parameters (retaining heads) to predict the CIS of each cache unit during chunked prefill and performs cache eviction based on the predicted CIS. We conduct extensive experiments across different models and multiple datasets to compare Locret with major efficient inference techniques, including offloading, quantization, eviction, and sparse attention. Results show that Locret outperforms all baselines, using less GPU memory and without requiring offloading to CPU memory. Furthermore, Locret can be combined with other efficient inference methods, such as KV cache quantization and attention pool-based token dropping. Locret maintains high inference throughput on consumer-grade devices and demonstrates robustness to variations in the intermediate size of retaining heads and training data. The complete framework of Locret, including both training and inference, highlights its suitability for low-resource computing scenarios. Locret can be applied to various application scenarios, such as end-side multi-modal model inference and context compression during disaggregated inference, which we plan to explore in future work. We also aim to test Locret on additional model architectures, including multi-latent models and models of varying sizes. Additionally, we are highly interested in integrating Locret into end-side deployment infra-structures to enhance inference efficiency, particularly for devices like smartphones and personal computers.
Acknowledgment
[…]
Appendix A Hyperparameters, Environment and Baselines
A.1 Training
A.2 Inference
Table 6: Hyperparameters in Locret’s inference stage.
Model $b$ $B$ $n_{s}$ $n_{loc}$
Phi-3-mini-128K 6000 3072 2500 100
Llama-3.1-8B-instruct 16384 1024 2500 100
A.3 System Environment
For all the experiments except the 4090 experiments in Section 4.3, we use a workstation with 8$\times$×Nvidia A800/H800 GPUs and 104 Intel(R) Xeon(R) Platinum 8470 CPUs. The device has 1.0 TB CPU memory. The operating system is Red Hat 4.8.5. We conduct all experiments except the full attention full KV cache inference on a single GPU, and 2 GPUs for full attention settings.
A.4 Baselines
We compare Locret with full attention inference, InfLLM, Quantization, SirLLM and MInference. FullAttn inference is performed using vllm (Kwon et al., 2023), which includes automatic tensor parallelism. InfLLM is a representative of the offloading-based methods, where the full KV cache is offloaded to CPU, and the most relavant blocks are retrieved to GPU during inference. For quantization method, we use the Hugging Face implementation of 2-bits KV cache quantization, which is inspired by Liu et al. (2024b), where quantization is conducted along channels instead of tokens. We denote this method as HF-2bits. SirLLM is an eviction-based token dropping algorithm, where tokens with low token-entropy is evicted once the cache is fullfilled. MInference is a typical method of reducing peak GPU memory consumption through rule-based sparse attention, but it does not reduce the size of KV cache. Note that InfLLM, HF-2bits and SirLLM does not have official implementation on Phi-3-mini-128K, thus we implement these three methods according to the original algorithm. We only use the short factor of RoPE for InfLLM, and no further model modification is conducted for HF-2bits and SirLLM.
Appendix B The Effect of Training
@@
Appendix C Orthogonality to Other Methods
C.1 Combination with Quantization

### v2

@@ 4.2 End-to-end Benchmark
4.3 Processing Speed on Real Consumer-Grade Devices
[…]
R.PassKey is a task where the model retrieves a 5-digit number from a large amount of irrelevant text, a task we believe to be relatively simple for humans. Thus, we consider the task to have failed if the accuracy falls below 95%. As shown in Table 2, aside from the settings that fail on this task, Locret achieves the highest inference speed among all methods that can correctly process R.PassKey:
(1) Due to its MHA architecture, Phi-3-mini-128K has a larger KV cache, which leads to failures for both HF-2bits and MInference. Storing the full KV cache on a single 4090 GPU is infeasible, as it requires 48GB of memory. Although the quantized KV cache is reduced to 6GB, converting representations requires significant GPU memory for its intermediate states, resulting in the failure of HF-2bits. While InfLLM can run in memory-limited scenarios, its offloading process slows down inference. SirLLM fails due to its inaccurate eviction, which cannot correctly identify the 5-digit number.
(2) In the GQA model (Llama-3.1-8B-instruct), which has a smaller KV cache, the quantized cache can fit within the GPU memory. However, the quantization and dequantization processes become the bottleneck, leading to significantly slower speed performance. The performance of InfLLM, SirLLM, and MInference is similar to that seen with Phi-3-mini-128K. Although MInference benefits from faster encoding speed, it fails this task because it cannot process the entire input sequence at once. Locret strikes a balance between inference speed and performance, making it a far more suitable solution for long-context scenarios on consumer-grade devices.
4.4 Locret-Q: Supporting Query-Driven Tasks
We mainly address question Q3 in this section. Query-driven tasks are characterized by highly sparse yet query-correlated critical regions within the context. A representative example is the Multikey-NIAH task from RULER (Hsieh et al., 2024), where the context comprises JSON key-value pairs given by UUIDs, and the objective is to retrieve specific values using given keys. Such contexts are inherently challenging to compress effectively without query information. As introduced by Sun et al. (2024), existing eviction-based techniques exhibit significant performance degradation when applied to query-driven tasks. Due to this, we evaluate Locret against selected eviction-based baselines using the RULER benchmark (Hsieh et al., 2024). We introduce Locret-Q, a query-aware variant of Locret. When training the retaining heads, we prepend the last $l_{q}$ query tokens to the sequence and gather CIS labels. At inference, the query is inserted at the sequence start, ensuring its visibility across all eviction actions. This adaptation enables Locret-Q to perform query-aware eviction.
Table 3 shows Locret-Q against SnapKV (Li et al., 2024a), H2O (Zhang et al., 2024d), and SirLLM (Yao et al., 2024) on RULER with a 128K context length. For reference, we also include results for FullAttn and MInference. Metrics include prefill speed, decode speed, and task performances are reported. As shown in Table 3, all efficient inference methods exhibit performance degradation to some extent. However, Locret-Q outperforms other eviction-based methods and even surpasses MInference, demonstrating its effectiveness on query-driven tasks. SnapKV shows performance degradation, while H2O and SirLLM fail completely on RULER. For speedup, methods combining eviction and chunked prefill (Locret-Q, Locret, and SirLLM) significantly reduce prefill time, achieving >2×>2$\times$> 2 × speedup over FullAttn. SnapKV cannot accelerate prefill due to no computation reduction. H2O suffers from extremely slow prefill as it relies on full-sequence attention scores, incompatible with efficient implementations like Flash-Attention. Decoding speeds are similar across eviction-based methods, as they maintain comparable KV cache sizes after eviction, and all of them are faster than FullAttn. Notably, Locret fails on RULER, showing a gap compared to Locret-Q, highlighting the necessity of query-awareness for query-centric tasks. A simple modification to Locret unlocks its potential for such tasks.
Table 3: Performance, prefill speed, and decode speed on RULER. The best and second-highest scores among eviction-based methods in each column are highlighted in bold and underlined. “” indicates only testing the first 20 entries per subtask due to poor performance. FullAttn is implemented using Flash-Attention.
Method RULER-128K (%) Prefill (tok/s) Decode (tok/s)
FullAttn 82.20 4319.95 12.32
MInference 72.97 7205.06 2.61
Locret-Q 75.54 9587.84 40.11
SnapKV 48.76 4203.34 36.49
H2O 15.04 464.73 44.70
SirLLM 13.23 9717.41 40.86
Locret 34.33 9587.09 37.38
4.5 Additional Experiments
Additional experiments that further investigate Locret are included in the appendices due to space limitations. We compare Locret with randomized eviction in Appendix E, evaluate Locret on LongBench in Appendix F, provide a hyperparameter analysis in Appendix G, and explore the combination of Locret with other methods in Appendix H. Please refer to the appendices for further details.
5 Conclusion
We propose Locret, a lightweight training-based method that enables memory-efficient long-context LLM inference on consumer-grade devices. Locret introduces retaining heads to predict the CIS of each cache unit during chunked prefill and performs accurate cache eviction. We conduct extensive experiments across different models and multiple datasets to compare Locret with major efficient inference techniques, and results show that Locret outperforms all baselines, using less GPU memory and without requiring offloading to CPU memory. Locret-Q, a query-aware variant of Locret, can further process query-centric tasks without significant performance degradation. Future work will involve testing Locret on other model architectures, such as encoder-decoder and multi-latent models. More evaluations of Locret on other popular devices like NVIDIA Jetson are also planned. We also plan to explore integrating existing KV cache budget allocation methods with Locret to further enhance inference efficiency.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Hyperparameters, Environment and Baselines
A.1 Training
A.2 Inference
Table 4: Hyperparameters in Locret’s inference stage. “$b$” is cache budget, “$B$” refers to chunk size of chunked prefill, “$n_{s}$” refers to stabilizers length and “$n_{loc}$” is local length.
Model $b$ $B$ $n_{s}$ $n_{loc}$
Phi-3-mini-128K 6000 3072 2500 100
Llama-3.1-8B-instruct 16384 1024 2500 100
A.3 System Environment
For all the experiments except the 4090 experiments in Section 4.3, we use a workstation with 8$\times$×NVIDIA A800/H800 GPUs and 104 Intel(R) Xeon(R) Platinum 8470 CPUs. We only use 1 GPU from the cluster for training, as the GPU requirements are less than 80GB for all training procedures. The device has 1.0 TB CPU memory. The operating system is Red Hat 4.8.5. We conduct all experiments except the full attention full KV cache inference on a single GPU, and 2 GPUs for full attention settings.
A.4 Baselines
We compare Locret with full attention inference, InfLLM, Quantization, SirLLM and MInference. FullAttn inference is performed using vllm (Kwon et al., 2023), which includes automatic tensor parallelism. InfLLM is a representative of the offloading-based methods, where the full KV cache is offloaded to CPU, and the most relavant blocks are retrieved to GPU during inference. For quantization method, we use the Hugging Face implementation of 2-bits KV cache quantization, which is inspired by Liu et al. (2024b), where quantization is conducted along channels instead of tokens. We denote this method as HF-2bits. SirLLM is an eviction-based token dropping algorithm, where tokens with low token-entropy is evicted once the cache is fullfilled. We use the official implementation of SirLLM, which includes some CPU operations including importance sorting. MInference is a typical method of reducing peak GPU memory consumption through rule-based sparse attention, but it does not reduce the size of KV cache. Note that InfLLM, HF-2bits and SirLLM does not have official implementation on Phi-3-mini-128K, thus we implement these three methods according to the original algorithm. We only use the short factor of RoPE for InfLLM, and no further model modification is conducted for HF-2bits and SirLLM.
Appendix B The Global and Local Discrepancy of Scoring Functions
Cache importance scoring functions can generally be categorized into two types: causal and non-causal. Non-causal functions, e.g. H2O and SnapKV, require information from subsequent cache units to determine the importance score of a cache unit, making them dependent on prefilling the entire sequence. On the other hand, causal functions, e.g. SirLLM and Locret, predict cache importance without relying on subsequent information. Non-causal scoring functions are incompatible with chunked prefill because they cannot calculate scores without access to the full sequence. If such functions are integrated with chunked prefill, they often face a significant discrepancy between the local importance score (without considering subsequent information) and the global importance score (with full context).
To investigate this discrepancy, we measure the consistency of the top 10% most important cache positions identified in prefixes of various lengths compared to the full context. For reference, the full context is truncated to 6K tokens. The results shown in Figure 1 highlights that scoring functions requiring future information, such as H2O and SnapKV, suffer from significant discrepancies when subsequent cache units are not considered. SirLLM, while also causal, shows notable inaccuracies, leading to performance degradation as demonstrated in Table 1 and Table 2.
[…]
Method R.Number E.Sum E.MC C.Debug Avg.$\uparrow$
FullAttn 97.12 17.92 55.46 23.10 48.40
H2O 3.39 15.35 45.41 20.57 21.18
SnapKV 2.54 15.44 41.92 21.43 20.33
Locret 97.46 16.82 46.29 29.71 47.57
Appendix C Pseudocode of Locret
We provide the pseudocode of Locretin this section, and we describe the inferece process of Locret in Algorithm 1.
Algorithm 1 Locret Inference
  Output:Generated tokens $x_{gen}$
  // Leave the last $n_{loc}$ out to make sure they are not evicted.
  chunk_positions $\leftarrow$ split_chunk(00, $x$.length() $-n_{loc}$, $B$)
  K_cache, V_cache, score_cache $\leftarrow$ [], [], []
  for chunk $\in$ chunk_positions do
    begin_pos, end_pos $\leftarrow$ chunk.begin_pos, chunk.end_pos
    K_chunk, V_chunk, score_chunk $\leftarrow$
[…]
    K_cache $\leftarrow$ Concat(K_cache, K_chunk)
    V_cache $\leftarrow$ Concat(V_cache, V_chunk)
    score_cache $\leftarrow$ Concat(score_cache, score_chunk)
    if chunk is not the last chunk then
       // Keep the last $n_{s}$ caches to maintain higher context continuity.
       score_cache[score_cache.length()-
       $n_{s}$:score_cache.length()] $\leftarrow$+$\infty$ + ∞
    end if
    indices $\leftarrow$ top-$b$(score_cache).indices
    K_cache, V_cache, score_cache = K_cache[indices],
    V_cache[indices], score_cache[indices]
  end for
  K_cache, V_cache, score_cache $\leftarrow$
  $\mathbf{M}$($x$[$x$.length()$-n_{loc}$:$x$.length()], K_cache, V_cache)
  $x_{gen}\leftarrow\mathbf{M}$.generate(K_cache, V_cache)
  return $x_{gen}$
Appendix D Evaluating Locret-Q on RULER
Table 6: Performance, prefill speed, and decode speed on RULER (Detailed). The best and second-highest scores among eviction-based methods in each column are highlighted in bold and underlined, respectively. “” indicates testing limited to the first 20 entries per subtask due to poor performance. FullAttn is implemented using Flash-Attention.
Method SG1 SG2 SG3 MK1 MK2 MK3 MV MQ VT CWE FWE QA1 QA2 Avg.
FullAttn 99.40 99.80 99.60 98.20 87.60 67.00 94.65 98.00 60.98 71.40 72.20 78.20 41.6 82.20
MInference 100.00 98.60 99.00 95.40 58.20 23.80 84.35 95.70 66.40 45.94 74.67 67.80 38.80 72.97
Locret-Q 100.00 99.80 99.60 75.00 98.80 98.60 66.95 85.50 52.64 30.90 80.27 53.20 40.80 75.54
SnapKV 100.00 82.80 11.60 70.40 6.20 1.40 68.50 78.40 49.04 31.90 50.07 51.20 32.40 48.76
H2O 20.00 0.00 0.00 0.00 0.00 0.00 3.75 2.50 1.00 0.00 53.33 100.00 15.00 15.04
SirLLM 0.00 5.00 5.00 0.00 5.00 5.00 6.25 8.75 12.00 0.00 80.00 30.00 15.00 13.23
Locret 100.00 45.00 35.00 10.00 5.00 0.00 20.00 17.50 69.00 46.50 73.33 20.00 5.00 34.33
To evaluate Locret-Q’s performance on query-centric tasks, we compare it with selected eviction-based baselines: SnapKV, H2O, SirLLM, and vanilla Locret. We also include FullAttn (implemented with Flash-Attention) and MInference for reference. The RULER benchmark consists of 500 synthetic queries per task, each with a context length of 128K tokens. All methods are tested on Llama-3.1-8B-instruct.
For Locret-Q and Locret, we set the budget size $b$ to 6000, chunk size $B$ to 4096, stabilizers length $n_{s}$ to 2500, and local length $n_{loc}$ to 100. For SnapKV, the voting window size is set to 100, with the last 100 tokens retained. For H2O, due to its reliance on full-sequence attention scores, we use a layer-wise chunked prefill pattern with a chunk size of 1024. Larger chunk size would result in an out-of-memory error. For SirLLM, we configure the start size to 4, recent size to 1000, and budget size to 6000. All evaluations are conducted on a single NVIDIA A800-80GB GPU.
For prefill and decode speed testing, all methods except H2O are implemented with Flash-Attention; H2O uses PyTorch’s vanilla attention due to its incompatibility with efficient attention implementations.
Appendix E Trained Retaining Heads vs. Random Eviction
[…]
Random 0.00 34.00 5.09 2.68 18.34 1.54 0.00 13.71 2.57 4.92
Locret 100.00 97.46 16.82 7.61 46.29 11.31 10.00 27.92 29.71 34.73
[…]
Appendix F Evaluation on LongBench
We conduct additional experiments to evaluate Locret on LongBench (Bai et al., 2024b), comparing it with baselines such as Full Attention, MInference, InfLLM, and SirLLM. For this evaluation, we used Phi-3-mini-128K with a retained head trained on LongAlign. To ensure a fair comparison, we excluded all Chinese subtasks from LongBench and focused solely on the English subtasks, as Phi-3-mini-128K was not specifically trained on Chinese corpora. The results are presented below. For Locret , we follow the hyperparameters presented in Table 4.
Table 8: LongBench scores of Locret compared with baselines.
Method gov_ report triviaqa narrative qa qmsum musique 2wikimqa multifield qa_en repobench -p qasper hotpotqa multi_ news trec passage_ retrieval_en passage _count samsum lcc Avg.$\uparrow$
FullAttn 33.35 86.38 18.21 19.51 19.82 33.37 49.82 58.02 41.07 43.06 26.57 67.00 93.50 2.97 23.15 51.86 41.73
MInference 32.94 86.87 19.46 19.57 18.85 33.30 49.14 58.98 40.31 43.56 26.35 68.00 89.00 2.10 25.58 53.68 41.73
SirLLM 32.92 85.61 21.08 21.59 24.32 34.97 48.52 59.15 40.17 47.00 26.44 65.50 63.00 3.00 23.11 51.83 40.51
InfLLM 25.96 84.87 20.83 19.61 13.63 27.43 41.29 55.73 30.51 38.05 25.36 64.50 10.00 7.50 0.28 61.59 32.95
Locret 33.46 82.39 24.56 23.35 25.12 35.93 52.77 57.16 40.17 48.70 26.41 62.00 83.00 3.00 26.37 52.61 42.31
We also report the maximum memory usage, including the GPU memory, the CPU memory, and the total maximum memory, alongside the average score on LongBench. For FullAttn, we exclude the maximum memory usage, aligning with Figure 4.
Table 9: Comparison of methods on LongBench and memory usage.
Method LongBench Max GPU Memory Max CPU Memory Total Max Memory
FullAttn 41.73 - - -
MInference 41.73 27.63 0.17 27.80
SirLLM 40.51 18.29 0.05 18.34
InfLLM 32.95 20.03 8.95 28.98
Locret 42.31 17.71 0.15 17.86
From the experiments above, Locret demonstrates the best overall performance and excels in the majority of subtasks. It outperforms all the baselines without any noticeable performance degradation while consuming less memory. Although MInference also avoids performance drops, it requires more GPU memory compared to Locret. SirLLM achieves comparable memory usage but shows some performance decline compared to FullAttn and Locret. InfLLM exhibits the most significant performance drop, and its offloading mechanism results in the highest CPU memory usage, making it the method with the largest total memory consumption. These results highlight Locret as an outstanding approach for evaluation on LongBench.
Appendix G Hyperparameter Analysis
We examine three key hyperparameters: budget, stabilizer length, and chunk size.
Budget To evaluate the robustness of Locret under different budget constraints, we compare the proposed method with SnapKV (Li et al., 2024a) using chunked prefill on LongBench (Bai et al., 2024b). As shown in Figure 5(a), when the budget size increases, Locret demonstrates a faster performance improvement compared to SnapKV.
Stabilizers Length As discussed in Figure 3, stabilizers play a crucial role in context retrieval tasks. However, in NLU tasks, the stability of $n_{s}$ remains relatively high. We evaluate QMSum with different stabilizer lengths $n_{s}$, with the budget set at 6000. As illustrated in Figure 5(b), performance remains consistent when $n_{s}$ is small. The observed performance degradation at larger $n_{s}$ values is due to the reduced space available for other cache units.
Chunk Size Executing long-context inference on hardware with varying GPU memory limitations choices of chunk size. When the chunk size changes, Locret shows stable performance. We test on the NQ dataset from L-Eval using multiple chunk sizes ranging from 256 to 4096. The results, shown in Figure 5(c), highlight the stability of $n_{s}$.
Figure 5: Scores of Locret under (a) various budgets; (b) various $n_{s}$; (c) various chunk size.
Appendix H Orthogonality to Other Methods
@@
KV cache quantization. According to Zhang et al. (2024c), eviction-based methods like H2O struggle with compatibility when combined with KV cache quantization. Quantization introduces significant disturbance in the estimation of heavy-hitters, leading to severe performance degradation. However, Locret is not affected by such issues and can be combined with quantization while maintaining most of its performance. Here, we compare the performance degradation caused by quantization on Locret with that of the full attention method using the same metrics. We use Quanto as the quantization backend and report the average L-Eval score with Llama-3.1-8B-instruct as the model backbone. Table H shows that the performance drop caused by quantization on Locret is only slightly higher than that observed with the full attention method, indicating that Locret is a quantization-friendly approach. More details of the experiment are provided in Appendix H.1.
@@
Token merging. As described in Section 2, token dropping can also be implemented through an attention pool. Attention pool-based methods (Xiao et al., 2024c; Cai et al., 2024a; Mu et al., 2024; Munkhdalai et al., 2024) merge adjacent tokens or cache units into an attention pool, maintaining a static cache size. These methods are orthogonal to Locret , as the evicted tokens can be merged into a small cache pool and retained in GPU memory. We conduct the following experiment to demonstrate that Locret can serve as an effective plug-in scoring function within such frameworks, enhancing performance without increasing memory budget. We select LoCoCo (Cai et al., 2024a) as a representative of the latest attention pool-based methods. LoCoCo maintains a cache set consisting of two parts: the heavy hitters and the convolved non-heavy hitters. During each chunked prefill step, LoCoCo first identifies a set of heavy hitters according to H2O (Zhang et al., 2024d), then applies 1-D convolution to the non-heavy hitters to compress them into a static size. By replacing H2O’s heavy-hitter scoring function with Locret, we retain the cache units with high CIS and convolve the others. We compare this combination with standalone LoCoCo and Locret on L-Eval using the Llama-3.1-8B-instruct backbone and report the average score across all selected tasks. As shown in Table H, Locret achieves a higher score than LoCoCo, and the combined algorithm outperforms both standalone methods. This suggests that Locret provides a more accurate scoring function compared to H2O, and the two methods complement each other, demonstrating their orthogonality. Further details of the experiment are provided in Appendix H.2.
@@
Method R.Number E.Sum E.MC C.Debug Avg.$\uparrow$
Locret 97.46 16.82 46.29 29.71 47.57
Locret+PyramidKV 99.66 15.82 48.03 30.00 48.38
H.1 Combination with Quantization